OpenStudy (anonymous):

@arjund12 I can solve to find the angle measure for A and the length of line BC..... all I need help with is finding AB and I can then do the rest.

OpenStudy (anonymous):

use the pythagorean thm

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (anonymous):

make the hypo x and us sin and solve for it

OpenStudy (anonymous):

sin theta = 30 /x x =30 sin(20)

OpenStudy (anonymous):

which will give you length of AB

OpenStudy (anonymous):

therefore, AB = 88 approximentally.

OpenStudy (anonymous):

I got 34, trying to find what I did wrong hold on

OpenStudy (phi):

do you use SOH CAH TOA to remember which trig function to use ? in a right triangle, once you know one angle (not including the 90 degree angle) and one side, you can find all sides and all angles.

OpenStudy (anonymous):

lostintime if possible show me your work, so i can tell you what you probably did wrong.

OpenStudy (phi):

you should probably use only trig

OpenStudy (anonymous):

dont use pythagorean thm

OpenStudy (anonymous):

use sin to find AB length

OpenStudy (anonymous):

Ohh okay so can you give me the equation. I'll do the work and then we can see if I do it right this time :]

OpenStudy (anonymous):

sin theta = opp/hypo in our case hypo = x

OpenStudy (anonymous):

theta = 20

OpenStudy (anonymous):

opp = 30

OpenStudy (anonymous):

thus, to find x we do x = 30/sin(20)

OpenStudy (anonymous):

which gives use 88 approximentally.

OpenStudy (anonymous):

yes I got it now! :] It's 88 cause you round the 87.7141 to and 88. Thanks soooo much! I can do the rest on my own now.

OpenStudy (anonymous):

ok, np.

OpenStudy (anonymous):

read my message i sent you.

OpenStudy (anonymous):

You do better teaching than my teacher does xD I seriously spents weeks on this and finally I understand how to do this :] Thanks so much again!

OpenStudy (anonymous):

*spent

OpenStudy (anonymous):

no problem.

OpenStudy (phi):

btw, if you want a good explanation , see https://www.khanacademy.org/math/trigonometry/basic-trigonometry/basic_trig_ratios/v/basic-trigonometry and the following video.

OpenStudy (anonymous):

Thanks for the link I'll save it :]